OpenStudy (anonymous):

@vtm386 Lessons.....

OpenStudy (anonymous):

ok im here

OpenStudy (anonymous):

OK lets start from factoring

OpenStudy (anonymous):

ok could you factor the problem \[3x^2 + 4x = 2\]

OpenStudy (anonymous):

I think no!

OpenStudy (anonymous):

For sure not, because 2 would be 2 times 1 and 2 &1 never add up to 4. See?

OpenStudy (anonymous):

ok so how would you do that problem?

OpenStudy (anonymous):

would you use the completing the square?

OpenStudy (anonymous):

At first,, in any case, I would subtract 2 from both sides.

OpenStudy (anonymous):

\[3x^2+4x=2\] \[3x^2+4x-2=2-2\] \[3x^2+4x-2=0\] this is the form, to even start choosing the method. Set it equal to zero. Right, @vtm386?

OpenStudy (anonymous):



OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

Now, lets see, have you ever factored anything before?